If $t$ is a real number, what is the maximum possible value of the expression $-t^2 + 8t -4$?
Answer by rothauserc(4718)

f(t) = -t^2 + 8t -4 is a parabola that curves downward
:
the t value for the vertex is -8 / -2 (-b/2a) = 4
:
now substitute t = 4 in f(t)
:
f(4) = -16 + 32 - 4 = 12
:
*************************************
the max possible value for f(t) is 12
*************************************
: